If the 12th term of an hp be 1/5 and the 19th term be 3/22 then the 4th term is 

Grade:9

2 Answers

Arun
4 years ago
Dear student
 
12th term of AP = 5
19th term of AP = 22/3
 
a + 11 d = 5
a + 18d = 22/3
Hence 7d = 7/3
d = 1/3
Hence a = 4/3
 
Hence 4th term of AP = a +3d = 7/3
 
Hence 4th term of HP = 3/7
Vikas TU
4 years ago
Dear student 
It is in HP 
So HP = 1/AP 
Solve like this , 
I think you get that , 
Hope this will help 
Good Luck